if the total charge stored is 240 micro Coulomb. Then calculate,
A, calculate the net potential difference
B, calculate the charge to be stored on C3
Given
C1=24 micro Coulomb
C2=4 micro Coulomb
C3=8 micro Coulomb
1
Expert's answer
2019-12-05T10:22:26-0500

Consider the following circuit:



First, we need to remember that charge is measured is coulombs and denoted as Q, while capacitance is measured in farads and denoted as C. We have four capacitors connected in parallel. The net potential difference measured by the voltmeter V equals the potential difference across each capacitor:


"V=\frac{Q_1}{C_1}=\frac{Q_2}{C_2}=\frac{Q_3}{C_3}=\frac{Q_4}{C_4}."

The net charge in a system of capacitors connected in parallel is


"Q=Q_1+Q_2+Q_3+Q_4,"

hence,


"Q_4=240-24-4-8=204 \space\mu\text{C}."

The net potential difference, as we see, depends on the capacitance "C_i" of each capacitor:


"C_1=\frac{Q_1}{V},C_2=\frac{Q_1}{V},C_3=\frac{Q_1}{V},C_4=\frac{Q_1}{V}."

Thus, we cannot calculate the net voltage because we do not know the capacitance of every capacitor. But we can calculate it if we substitute each capacitance in microfarads ("\mu \text{F}") to the first equation.